# Grandmaster-Level Chess Without Search

### v1

@@ 3 Results
Here we present our comprehensive experimental evaluation. For all parameters not explicitly mentioned we use the same setting across our two main experiments (Section 3.1, Section 3.2); for investigating scaling behavior and all ablations in Section 3.3 and Section 3.4 we use a different set of default settings (geared towards getting representative results with better computational efficiency). We provide all details in Section A.2 and Section A.3, respectively.
@@ 3.1 Main Result
In Table 1 we show the playing strength (internal tournament Elo, external Lichess Elo, and puzzle solving competence) and predictor metrics of our large-scale transformer models when trained on the full (M games) training set. Our main evaluation compares three transformer models with M, M, and M parameters after training (none of them overfit the training set as shown in Section B.1). The results show that all three models exhibit non-trivial generalization to novel boards and can successfully solve a large fraction of puzzles. Across all metrics, having larger models consistently improves scores, confirming that model scale matters for strong chess performance. Our largest model achieves a blitz Elo of 2895 against human players, which places it into grandmaster territory. However, the Elo drops when playing against bots on Lichess, which may be a result of having a significantly different player-pool, some minor technical issues, and perhaps a qualitative difference in how bots exploit weaknesses compared to humans (see Section 5 for a detailed discussion of these issues).
@@ 3.2 Puzzles
In Figure 2 we compare the puzzle performance of our 270M parameter model against Stockfish 16 (time limit of 50ms per move), GPT-3.5-turbo-instruct, and AlphaZero’s value network. We use our large puzzle set of k puzzles, grouped by their assigned Elo difficulty from Lichess. Stockfish 16 performs the best across all difficulty categories, followed by our 270M model. AlphaZero’s value network (trained on M games) and GPT-3.5-turbo-instruct achieve non-trivial puzzle performance, but significantly lag behind our model. We emphasize that solving the puzzles requires a correct move sequence, and since our policy cannot explicitly plan ahead, solving the puzzle sequences relies entirely on having good value estimates that can be used greedily.
3.3 Scaling “Laws”
Figure 3 shows our scaling analysis over the dataset and model size. We visualize the puzzle accuracy (training and test loss in Figure A4), which correlates well with the other metrics and the overall playing strength. For small training set size (k games, left panel) larger architectures (M) start to overfit as training progresses. This effect disappears as the dataset size is increased to k (middle panel) and M games (right panel). The results also show that the final accuracy of a model increases as the dataset size is increased (consistently across model sizes). Similarly, we observe the general trend of increased architecture size leading to increased overall performance regardless of dataset size (as also shown in our main result in Section 3.1).
@@ 3.4 Variants and Ablations
Predictor-target | AV | 83.3 | 63.0 | 0.259 |
SV | 77.5 | 58.5 | 0.215 | |
BC | 65.7 | 56.7 | 0.116 | |
Network depth | 2 | 62.3 | 54.4 | 0.219 |
4 | 76.2 | 59.9 | 0.242 | |
8 | 81.3 | 62.3 | 0.254 | |
16 | 80.4 | 62.3 | 0.255 | |
Data sampler | Uniform | 83.3 | 63.0 | 0.259 |
Weighted | 49.9 | 48.2 | 0.192 | |
Value bins | 16 | 83.0 | 61.4 | 0.248 |
@@
256 | 83.7 | 63.0 | 0.260 | |
Loss function | (class.) | 81.3 | 62.3 | 0.254 |
L2 (regr.) | 82.6 | 58.9 | 0.235 | |
Stockfish Limit [s] | 0.05 | 84.0 | 62.2 | 0.256 |
0.1 | 85.4 | 62.5 | 0.254 | |
0.2 | 84.3 | 62.6 | 0.259 | |
0.5 | 83.3 | 63.0 | 0.259 |
We test a series of experimental variants and perform extensive ablations using the M parameter model. The results and conclusions drawn are used to inform and justify our design choices and determine default model-, data-, and training-configurations. Table 2 summarizes all results.
Predictor-targets
By default we learn to predict action-values given a board state. Here we compare against using state-values or oracle actions (behavioral cloning) as the prediction targets. See Section 2.3 and Figure 1 for more details and how to construct policies from each of the predictors. As the results in Table 2 show, the action-value predictor is superior in terms of action-ranking (Kendall’s ), action accuracy, and puzzle accuracy. The same trend is also shown in Figure A5 (in Section B.2, which tracks puzzle accuracy over training iterations for the different predictors. This superior performance of action-value prediction might stem primarily from the significantly larger action-value dataset (B state-action pairs vs. M states for our largest training set constructed from M games). We thus run an additional ablation where we train all three predictors on exactly the same amount of data—results shown in Section B.2 largely confirm this hypothesis. Please see our more detailed discussion of the different predictor targets as we discuss these results in Section B.2, where we also discuss performance discrepancy between behavioral cloning and the state-value predictor based policy may be largely explained by the fact that we train on expert’s actions only instead of the full action distribution of the expert.
@@
We show the influence of increasing the transformer’s depth while keeping the number of parameters constant in Table 2. Since transformers may learn to roll out iterative computation (which arises in search) across layers, deeper networks may hold the potential for deeper unrolls. We compensate for having fewer layers by varying the embedding dimension and widening factor such that all models have the same number of parameters. The performance of our models increases with their depth but seems to saturate at around 8 layers, indicating that depth is important, but not beyond a certain point.
Data sampler
We remove duplicate board states during the generation of the training and test sets. This increases data diversity but introduces distributional shift compared to the “natural” game distribution of boards where early board states and popular openings occur more frequently. To quantify the effect of this shift we use an alternative “weighted” data sampler that draws boards from our filtered training set according to the distribution that would occur if we had not removed duplicates. Results in Table 2 reveal that training on the natural distribution (via the weighted sampler) leads to significantly worse results compared to sampling uniformly randomly from the filtered training set (both trained models are evaluated on a filtered test set with uniform sampling, and the puzzle test set). We hypothesize that the increased performance is due to the increased data diversity seen under uniform sampling. As we train for very few epochs, the starting position and common opening positions are only seen a handful of times during training under uniform sampling, making it unlikely that strong early-game play of our models can be attributed to memorization.
@@
Table 2 shows the impact of varying the number of bins used for state- and action-value discretization (from to ), demonstrating that more bins lead to improved performance. To strike a balance between performance and computational efficiency, we use bins for our ablations and for the main experiments.
Loss function
We treat learning Stockfish action-values as a classification problem and thus train by minimizing cross-entropy loss (log-loss). This is as close as possible to the (tried and tested) standard LLM setup. An alternative is to treat the problem as a scalar regression problem. If we parameterize a fixed-variance Gaussian likelihood model with a transformer and perform maximum (log) likelihood estimation, this is equivalent to minimizing mean-squared error (L2 loss). To that end, we modify the architecture to output a scalar (without a final log-layer or similar). The log-loss outperforms the L2 loss on two out of the three metrics (Table 2).
Stockfish time limit
We create training sets from million games annotated by Stockfish with varying time limits to manipulate the playing strength of our oracle. We report scores on the puzzle set (same for all models) and a test set created using the same time limit as the training set (different for all models). Table 2 shows that a basic time-limit of seconds gives only marginally worse puzzle performance. As a compromise between computational effort and final model performance we thus choose this as our default value (for our M games dataset we need about B action-evaluation calls with Stockfish, i.e., roughly 8680 days of unparallelized Stockfish evaluation time).
@@ 4 Related Work
Early chess AI research made heavy use of designing explicit search strategies coupled with heuristics, as evidenced by Turing’s initial explorations (Burt, 1955) and implementations like NeuroChess (Thrun, 1994). This approach culminated in systems like Deep Blue (Campbell et al., 2002) and Stockfish (Romstad et al., 2008), known for their advanced search algorithms. The development of AlphaZero (Silver et al., 2017) marked a paradigm shift, employing deep RL with Monte Carlo Tree Search, thus learning its own heuristics (policy and value networks) instead of manually designing them. Neural networks play a significant role in chess AI (Klein, 2022), including enhancements to AlphaZero’s self-play mechanisms (V. et al., 2018), the use of deep RL (Lai, 2015), and a general trend of moving away from explicit search methods, by leveraging large-scale game datasets for training (David et al., 2016; Schrittwieser et al., 2020).
The rise of large language models has also led to innovations in chess AI, cf. Kamlish’s language-based models (Kamlish et al., 2019), the encoding of chess games via natural language (Toshniwal et al., 2022; DeLeo and Guven, 2022), and the evaluation LLMs ability to play chess (Carlini, 2023; Gramaje, 2023). Czech et al. (2023) show that strategic input representations and value loss enhancements significantly boost chess performance of vision transformers, and Alrdahi and Batista-Navarro (2023); Feng et al. (2023) show that adding chess specific data sources (e.g., chess textbooks) to language model training can improve their chess performance. Stöckl (2021) explored scaling effects of transformers on chess performance, which resonates with our emphasis on the importance of model and dataset scale.
5 Discussion
In order to use our state-based policies to play against humans and bots, two minor technical issues appear that can only be solved by having (some) access to game history. We briefly discuss both issues and present our workarounds.
Blindness to threefold repetition
By construction, our state-based predictor cannot detect the risk of threefold repetition (drawing because the same board occurs three times), since it has no access to the game history (FENs contain minimal historical info, sufficient for the Fifty Move rule). To reduce draws from threefold repetitions, we check if the bot’s next move would trigger the rule and set the corresponding action’s win percentage to before computing the softmax. However, our bots still cannot plan ahead to minimize the risk of being forced into threefold repetition.
Indecisiveness in the face of overwhelming victory
If Stockfish detects a mate-in- (e.g., or ) it outputs and not a centipawn score. We map all such outputs to the maximal value bin (i.e., a win percentage of ). Similarly, in a very strong position, several actions may end up in the maximum value bin. Thus, across time-steps this can lead to our agent playing somewhat randomly, rather than committing to one plan that finishes the game quickly (the agent has no knowledge of its past moves). This creates the paradoxical situation that our bot, despite being in a position of overwhelming win percentage, fails to take the (virtually) guaranteed win and might draw or even end up losing since small chances of a mistake accumulate with longer games (see Figure 4). To prevent some of these situations, we check whether the predicted scores for all top five moves lie above a win percentage of and double-check this condition with Stockfish, and if so, use Stockfish’s top move (out of these) to have consistency in strategy across time-steps.
Elo: Humans vs. bots
Table 1 shows a difference in Lichess Elo when playing against humans compared to bots. While the precise reasons are not entirely clear, we have three plausible hypotheses: (i) humans tend to resign when our bot has overwhelming win percentage but many bots do not (meaning that the previously described problem gets amplified when playing against bots); (ii) humans on Lichess rarely play against bots, meaning that the two player pools (humans and bots) are hard to compare and Elo ratings between pools may be miscalibrated (Justaz, 2023); and (iii) based on preliminary (but thorough) anecdotal analysis by a chess NM, our models make the occasional tactical mistake which may be penalized qualitatively differently (and more severely) by other bots compared to humans (see some of this analysis in Sections B.4 and B.5). While investigating this Elo discrepancy further is interesting, it is not central to our paper and does not impact our main claims.
5.1 Limitations
While our largest model achieves very good performance, it does not completely close the gap to Stockfish 16. All our scaling experiments point towards closing this gap eventually with a large enough model trained on enough data. However, the current results do not allow us to claim that the gap can certainly be closed. Another limitation, as discussed earlier, is that our predictors see the current state but not the complete game history. This leads to some fundamental technical limitations that cannot be overcome without small domain-specific heuristics or augmenting the training data and observable info. Finally, when using a state-value predictor to construct a policy, we consider all possible subsequent states that are reachable via legal actions. This requires having a transition model , and may be considered a version of 1-step search. While the main point is that our predictors do not explicitly search over action sequences, we limit the claim of ‘without search’ to our action-value policy and behavioral cloning policy.
Note that the primary goal of this project was to investigate whether a complex, search-based algorithm, such as Stockfish 16, can be well approximated with a feedforward neural network. In the course of this, we have made a serious attempt to produce a strong chess policy and estimate its playing strength, but we have not exhausted every conceivable option to maximize playing strength—it may well be that further tweaks of our approach could lead to even stronger policies. Similarly, we have made a serious attempt at calibrating our policy’s playing strength via Lichess, where the claim of “grandmaster-level” play currently holds against human opponents, but we have not calibrated our policy under official tournament conditions. We also cannot rule out that opponents, through extensive repeated play, may be able to find and exploit weaknesses reliably due to the fairly deterministic nature of our policy.
@@ 6 Conclusion
Our paper shows that is is possible to distill an approximation of Stockfish 16 into a feed-forward transformer via standard supervised training. The resulting predictor generalizes well to unseen board states, and, when used in a policy, leads to strong chess play (Lichess Elo of 2895 against humans). We demonstrate that strong chess capabilities from supervised learning only emerge at sufficient dataset and model scale. Our work thus adds to a rapidly growing body of literature showing that complex and sophisticated algorithms can be distilled into feed-forward transformers, implying a paradigm-shift away from viewing large transformers as “mere” statistical pattern recognizers to viewing them as a powerful technique for general algorithm approximation.
Impact Statement
While the results of training transformer-based architectures at scale in a (self) supervised way will have significant societal consequences in the near future, these concerns do not apply to a closed domain like chess that has limited real-world impact and has been a domain of machine superiority for decades. Another advantage of supervised training on a single task over other forms of training (particularly self-play or reinforcement learning, and meta-learning) is that the method requires a strong oracle solution to begin with (for data annotation) and is unlikely to significantly outperform the oracle—so the potential for the method to rapidly introduce substantial unknown capabilities (with wide societal impacts) is very limited.

### v2

@@ 3 Case-Study Results
We use two settings for our experiments: (i) a large-scale setting for our main results (Sections 3.1 and 3.2; details in Section A.2), and (ii) a setting geared towards getting representative results with better computational efficiency for our ablations (Sections 3.3 and 3.4; details in Section A.3).
@@ 3.1 Main Result
Lichess Elo | ||||||||
Agent | Train | Search | Input | Tournament Elo | vs. Bots | vs. Humans | Puzzle Acc. (%) | |
9M Transformer (ours) | SL | FEN | 2025 () | 2054 | - | 88.9 | ||
136M Transformer (ours) | SL | FEN | 2259 () | 2156 | - | 94.5 | ||
270M Transformer (ours) | SL | FEN | 2299 () | 2299 | 2895 | 95.4 | ||
GPT-3.5-turbo-instruct | SSL | PGN | - | 1755 | - | 66.5 | ||
AlphaZero (policy net only) | RL | PGN | 1777 () | - | - | 56.1 | ||
AlphaZero (value net only) | RL | PGN | 1992 () | - | - | 82.0 | ||
AlphaZero ( MCTS sim.) | RL | ✓ | PGN | 2470 () | - | - | 95.6 | |
Leela Chess Zero (policy net only) | RL | PGN | 2292 () | 2224 | - | 88.6 | ||
Leela Chess Zero (value net only) | RL | PGN | 2418 () | 2318 | - | 95.9 | ||
Leela Chess Zero ( MCTS sim.) | RL | ✓ | PGN | 2858 () | 2620 | - | 99.6 | |
Stockfish 16 (ms per move) [oracle] | SL | ✓ | FEN + Moves | 2711 () | 2713 | - | 99.8 | |
Stockfish 16 (s per board) | SL | ✓ | FEN + Moves | 2935 ( | 2940 | - | 100.0 |
Table 1 shows the playing strength (internal tournament Elo, external Lichess Elo, and puzzle accuracy) of our large-scale transformers trained on the full (M games) training set. We compare models with M, M, and M parameters (none of them overfit the training set as shown in Section B.4). The results show that all three models exhibit non-trivial generalization to novel boards and can successfully solve a large fraction of puzzles. Across all metrics, increasing model size consistently improves the scores, confirming that model scale matters for strong chess performance. Our largest model achieves a blitz Elo of 2895 against human players, which places it into grandmaster territory. However, the Elo drops when playing against bots on Lichess, which may be a result of having a significantly different player pool [25], minor technical issues, or a qualitative difference in how bots exploit weaknesses compared to humans (see Section 4 for a detailed discussion).
@@ 3.2 Puzzles
In Figure 2 we compare the puzzle performance of our 270M parameter model against Stockfish 16 (50ms limit per move), GPT-3.5-turbo-instruct, AlphaZero, and Leela Chess Zero. We use our large puzzle set of K puzzles, grouped by their assigned Elo difficulty from Lichess. Stockfish 16 performs the best across all difficulty categories, followed by Leela Chess Zero, AlphaZero, and our model. Impressively, our model, which does not use any explicit search at test time, nearly matches the performance of AlphaZero with (Monte Carlo) tree search. GPT-3.5-turbo-instruct achieves non-trivial puzzle performance but significantly lags behind our model. We emphasize that solving the puzzles requires a correct move sequence, and since our policy cannot explicitly plan ahead, solving the puzzle sequences relies entirely on having good value estimates that can be used greedily.
3.3 Scaling Analysis
Figure 3 shows a scaling analysis over the dataset and model size. We visualize the puzzle accuracy (train and test losses in Figure A3), which correlates well with the other metrics and the overall playing strength. For small training set size (K games, left panel) larger architectures (M) start to overfit. This effect disappears as the dataset size is increased to K (middle panel) and M games (right panel). The results also show that the final accuracy of a model increases as the dataset size is increased (consistently across model sizes). Similarly, we observe the general trend of increased architecture size leading to increased overall performance (as in our main result in Section 3.1).
@@ 3.4 Variants and Ablations
We perform extensive ablations using the M-parameter model and show the results in Tables 2 and A2. We use the results and conclusions drawn to inform and justify our design choices and determine the default model-, data-, and training configurations.
@@
Predictor target | Action-Value | 83.3 | 63.0 | 0.259 |
State-Value | 77.5 | 58.5 | 0.215 | |
Behavioral Cloning | 65.7 | 56.7 | 0.116 | |
Loss function | HL-Gauss (class.) | 82.0 | 61.8 | 0.257 |
(class.) | 80.6 | 61.9 | 0.257 | |
L2 (regr.) | 80.8 | 58.9 | 0.240 | |
Network depth | 2 | 54.7 | 51.5 | 0.209 |
4 | 40.5 | 44.3 | 0.179 | |
8 | 79.5 | 60.7 | 0.252 | |
16 | 81.3 | 61.6 | 0.256 | |
32 | 79.5 | 61.2 | 0.254 | |
No. of value bins | 16 | 83.0 | 61.4 | 0.248 |
@@
256 | 83.7 | 63.0 | 0.260 |
Predictor targets
By default we learn to predict action-values given a board state. Here we compare against using state-values or oracle actions (behavioral cloning) as the prediction targets (Sections 2.3 and 1 describe how to construct policies from the predictors). Tables 2 and A4 show that the action-value predictor is superior in terms of action-ranking (Kendall’s ), action accuracy, and puzzle accuracy. This superior performance might stem primarily from the significantly larger action-value dataset (B state-action pairs vs. M states for our largest training set). We thus run an ablation where we train all three predictors on the same amount of data (results in Tables A4 and A5, which largely confirm this hypothesis). Section B.5 contains detailed discussion, including why the performance discrepancy between behavioral cloning and the state-value prediction policy may be largely due to training only on expert actions rather than the full action distribution.
Loss function
We treat learning Stockfish action-values as a classification problem and train by minimizing the HL-Gauss loss [18]. Here we compare this to the cross-entropy loss (log-loss), which is as close as possible to the (tried and tested) standard LLM setup. Another alternative is to treat the problem as a scalar regression problem by parameterizing a fixed-variance Gaussian likelihood model with a transformer and performing maximum (log) likelihood estimation, i.e., minimizing the mean-squared error (L2 loss). To that end, we modify the architecture to output a scalar (without a final log-layer or similar). Table 2 shows that the HL-Gauss loss outperforms the other two losses.
@@
Table 2 shows the influence of increasing the transformer’s depth (i.e., the number of layers) for a fixed number of parameters (we vary the embedding dimension and the widening factor such that all models have the same number of parameters). Since transformers may learn to roll out iterative computation (which arises in search) across layers, deeper networks may hold the potential for deeper unrolls. The performance of our models increases with their depth but saturates at around 16 layers, indicating that depth is important, but not beyond a certain point.
@@
Table 2 shows the impact of varying the number of bins used for state- and action-value discretization (from to ), demonstrating that more bins generally lead to improved performance (up to a certain point). We therefore use bins for our experiments.
4 Discussion
When using our state-based policies to play against humans and bots, two minor technical issues appear that can only be solved by having (some) access to the game’s history. In this section, we discuss both issues and present our corresponding workarounds.
Blindness to threefold repetition
By construction, our state-based predictor cannot detect the risk of threefold repetition (drawing because the same board occurs three times), since it has no access to the game’s history (FENs contain minimal historical info, sufficient for the fifty-move rule). To reduce draws from threefold repetitions, we check if the agent’s next move would trigger the rule and set the corresponding action’s win percentage to before computing the . However, our agents still cannot plan ahead to mimize the risk of being forced into threefold repetition.
Indecisiveness in the face of overwhelming victory
If Stockfish detects a mate-in- it outputs and not a centipawn score. When annotating our dataset, we map all such outputs to a win percentage of . Similarly, in a very strong position, several actions may end up in the maximum value bin. Thus our agent sometimes plays somewhat randomly rather than committing to a plan that finishes the game (the agent has no knowledge of its past moves). Paradoxically, this means that our agent, despite being in a position of overwhelming win percentage, sometimes fails to take the (virtually) guaranteed win (see Figure 4) and might draw or even end up losing since small chances of a mistake accumulate with longer games. To alleviate this, we check whether the predicted scores for all top five moves lie above a win percentage of and double-check this condition with Stockfish, and if so, use Stockfish’s top move (out of these) to have consistency in strategy across time-steps.
Elo: Humans vs. bots
We have three plausible hypotheses for why the Lichess Elo in Table 1 differs when playing against humans vs. bots: (i) humans tend to resign when our bot is in an overwhelmingly winning position but bots do not (i.e., the previously described problem gets amplified against bots); (ii) most humans on Lichess rarely play against bots, i.e., the two player pools (humans and bots) are hard to compare and their Elo ratings may be miscalibrated [25]; and (iii) based on anecdotal analysis by a chess National Master, our models make the occasional tactical mistake which may be penalized more severely by bots than humans (analysis in Sections B.7 and B.8). While investigating this Elo discrepancy is interesting, it is not central to our paper and does not impact our main claims.
4.1 Limitations
Our primary goal was to investigate whether a complex search algorithm such as Stockfish 16 can be approximated with a feedforward neural network on our dataset via supervised learning. While our largest model achieves good performance, it does not fully close the gap to Stockfish 16, and it is unclear whether further scaling would close this gap or whether other innovations are needed.
While we produced a strong chess policy, our goal was not to build a state-of-the-art chess engine. Our models are impractical in terms of speed and would perform poorly in computer chess tournaments with compute limitations. Therefore, we calibrated our policy’s playing strength via Lichess, where the claim of “grandmaster-level” play currently holds only against human opponents. In addition, we only evaluated our biggest model against humans on Lichess due to the extensive amount of time required. We also cannot rule out that opponents, through extensive repeated play, may find weaknesses due to the deterministic nature of our policy. Finally, we reiterate that we compare to other engines to situate our models within the wider landscape, but that a direct comparison between all the engines comes with a lot of caveats due to differences in their inputs (FENs vs. PGNs), training protocols (RL vs. supervised learning), and the use of search at test time.
Leela Chess Zero’s networks, which are trained with self-play and RL, achieve higher Elo ratings without using explicit search at test time than our transformers, which we trained via supervised learning. However, in contrast to our work, very strong chess performance (at low computational cost) is the explicit goal of this open source project (which they have clearly achieved via domain-specific adaptations). We refer interested readers to [6] (which was published concurrently to our work) for details on the current state-of-the-art and a comparison against our networks.
@@ 5 Related Work
Unsurprisingly given its long history in AI, there is a vast literature on applying algorithmic techniques to chess. Earlier works predominantly focused on methods which would increase the strength of chess playing entities, but as time has progressed, the role of chess as a problem domain has changed more from a challenge area to that of an illuminating benchmark, and our work continues this tradition.
Early computer chess research focused on designing explicit search strategies coupled with heuristics, as evidenced by Turing’s initial explorations [26] and implementations like NeuroChess [27]. This culminated in systems like Deep Blue [28] and early versions of Stockfish [29].
AlphaZero [7] and Leela Chess Zero [5] marked a paradigm shift: They employed deep RL with Monte Carlo tree search to learn heuristics (policy and value networks) instead of manually designing them [30, 31]. Several works built upon this framework [32], including enhancements to AlphaZero’s self-play mechanisms [33] and the use of model-free RL [23]. At the time of writing, Leela Zero’s T82 networks were state-of-the art, but their training protocol, training data, and some architectural details were not fully reported. In the meantime, Monroe and Leela Chess Zero Team [6] have published a detailed tech report on another transformer-based Leela Zero architecture called the ChessFormer. Their comparison shows that ChessFormers comparable in size to our models outperform our vanilla transformers while requiring fewer FLOPS thanks to clever domain-specific adaptations.
Another line of work moved away from explicit search methods by leveraging large-scale game datasets for (un)supervised learning, both for chess [34, 35, 36, 37] and Go [38]. Most closely related to our work, project Maia [35] used behavioral cloning on 12M human games but, rather than maximizing performance (i.e., distilling Stockfish), focused on predicting human moves at different levels of play.
The rise of large (pretrained) foundation models also led to innovations in various areas of computer chess research: learning the rules of chess [39, 40], evaluating move quality [41], evaluating playing strength [9, 42], state tracking [43, 44], and playing chess from visual inputs [45]. Fine-tuning on chess-specific data sources (e.g., chess textbooks) has further improved performance [46, 47, 48].
@@ 6 Conclusion
Our paper introduces ChessBench, a large-scale, open source benchmark dataset for chess, and shows the feasibility of distilling an approximation of Stockfish 16, a complex planning algorithm, into a feed-forward transformer via standard supervised training. The resulting predictor generalizes well to unseen board states, and, when used in a policy, leads to strong chess play. We show that strong chess capabilities from supervised learning only emerge at sufficient dataset and model scale. Our work thus adds to a rapidly growing body of literature showing that sophisticated algorithms can be distilled into feed-forward transformers, implying a paradigm shift to viewing large transformers as a powerful technique for general algorithm approximation rather than “mere” statistical pattern recognizers. Nevertheless, perfect distillation of Stockfish 16 is still beyond reach and closing the performance gap might need other (e.g., architectural) innovations. Our open source benchmark dataset, ChessBench, thus provides a solid basis for scientific comparison of any such developments.
7 Impact Statement
While the results of training transformer-based architectures at scale in a (self-)supervised way will have significant societal consequences in the near future, these concerns do not apply to closed domains, such as chess, that have limited real-world impact. Moreover, chess has been a domain of machine superiority for decades. Another advantage of supervised training on a single task over other forms of training (particularly self-play or reinforcement learning and meta-learning) is that the method requires a strong oracle solution to begin with (for data annotation) and is unlikely to significantly outperform the oracle, which means that the potential for the method to rapidly introduce substantial unknown capabilities (with wide societal impacts) is very limited.